OpenStudy (anonymous):

Integrate (-2x-2)e^x

OpenStudy (bahrom7893):

split this into two integrals: -2*Integral(x*e^x)-2*Integral(e^x) for the first integral use integration by parts.

OpenStudy (anonymous):

That integration by parts! That's what wolfram suggested...how does it work?

OpenStudy (anonymous):

:(

OpenStudy (bahrom7893):

let u=x, du=dx, dv=e^xdx, v=e^x

OpenStudy (bahrom7893):

The integral becomes: u*v-Int(v*du)=x*e^x-Integral(e^x*dx)

OpenStudy (anonymous):

ok thanks.

OpenStudy (amistre64):

you can int by parts without expanding first as well .. but potato potato ...